BOB & MARY EACH HAVE APPLES
IF BOB WERE TO GIVE MARY 2 OF HIS APPLES MARY WOULD HAVE TWICE AS MANY APPLES AS BOB. HOWEVER, IF MARY WERE TO GIVE BOB 2 OF HER APPLES, BOTH MARY AND BOB WOULD HAVE THE SAME NUMBER OF APPLES. HOW MANY APPLES DOES BOB AND MARY TO START?

Answer by fractalier(6550)

Let b and m be what Bob and Mary have before giving any of them away.
The first statement says
2(b - 2) = m + 2
The second says
m - 2 = b + 2
Solve both for m and set them equal...we have
2b - 4 = m + 2
m = 2b - 6
and
m = b + 4
so that
2b - 6 = b + 4
and you can find that
b = 10
and
m = 14
and you can check that it is true by plugging these numbers into the problem...